OpenStudy (anonymous):

Which exponetial function goes through the points 1,8 and 4,512 f(x)=4(2)^x f(x)=8(2)^-x f(x)=4(4)^-x f(x)=2(4)^x

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

D, f(x) = 2(4)^x x = 1 x = 4 f(1) = 2(4)^1 f(1) = 2(4) f(1) = 8 ----> (1,8) f(4) = 2(4)^4 f(4) = 2(256) f(4) = 512 ------> (4,512)

OpenStudy (anonymous):

your the best!!!!!!!!!!

OpenStudy (the_fizicx99):

np

OpenStudy (anonymous):

can you help with 1 more

OpenStudy (anonymous):

@tHe_FiZiCx99